OpenStudy (anonymous):

Claire purchases a new dress for the prom. The dress is priced $160, but it is on sale for 30% off. Claire’s aunt works at the store and can give her an additional 10% off. If the sales tax is 7.5%, how much does Claire pay for the dress? a. $93.24 b. $103.20 c. $108.36 d. $120.40

OpenStudy (anonymous):

10% of 160 is 16. 16 x 4=64 tax:160x7.5= 1200/100=$12 64-12=52 160-52=$108

OpenStudy (anonymous):

im probably wrong but good luck ;)